OpenStudy (anonymous):

|2x-5|=6 I know the first answer is 11/5, but how do i find the second answer

OpenStudy (dangerousjesse):

|x|=|-x|

OpenStudy (anonymous):

\[\left| 2x-5 \right|=6\] \[2x-5=\pm 6\]

jhonyy9 (jhonyy9):

how you get 11/5 ?

OpenStudy (anonymous):

solve 2x-5=6,find x and 2 x-5=-6,find x